OpenStudy (anonymous):

A sequence{an} is given by a1=sqrt(2), an+1=sqrt(2+an). a) by induction or otherwise, show that {an} is increasing and bounded above by 3. Apply the Monotonic Sequence Theorem to show that lim n-->infinity an exists. b) Find lim n-->infinity an.

OpenStudy (anonymous):

\[a_{n+1}=\sqrt{a_n+2}\]

OpenStudy (anonymous):

a) in order for \(a_n\) to be increasing we must show that \(a_{n+1}>a_n\) for \(n\ge1\)

OpenStudy (anonymous):

u r offline so i go to give a complete solution

mathslover (mathslover):

No need to give complete soln @mukushla

OpenStudy (anonymous):

first show that \(a_n<2\) by mathematical induction : Proof: For each positive integer \(n\ge1\) , let \(S(n)\) be the statement \(a_n<2\) Basis step: \(S(1)\) is the statement \(\sqrt{2}<2\) . Thus \(S(1)\) is true. Inductive step: We suppose that \(S(k)\) is true and prove that \(S(k+1)\) is true. Thus, we assume that \(a_k<2\) and prove that \(a_{k+1}<2\) we have \(a_{k+1}=\sqrt{a_k+2}<\sqrt{2+2}=2\) hence proved

mathslover (mathslover):

You did it :(

OpenStudy (anonymous):

o sorry

OpenStudy (anonymous):

i think thats against the code of contact...ok for other parts i will just give some hints

OpenStudy (anonymous):

in order to showing the sequence is increasing :\[a_{n+1}-a_n=\sqrt{a_n+2}-a_n=\frac{(\sqrt{a_n+2}-a_n)(\sqrt{a_n+2}+a_n)}{(\sqrt{a_n+2}+a_n)}=\frac{a_n+2-a_n^2}{(\sqrt{a_n+2}+a_n)}\]show that the later expression>0 using the result of induction

OpenStudy (anonymous):

i think u got the other parts from here :)